Sharp Industries has current assets of $10,000, shareholders’ equity of $27,000, current liabilities of $4,000, and long-term liabilities of $8,000. How much are the long-term assets?
Answer #1

Total liabilities=Current liabilities+long term liabilities

=(4000+8000)=$12000

Total assets=Total liabilities+Total equity

=(12000+27000)=$39000

Total assets=current assets+long term assets

long term assets=(39000-10,000)

=$29,000